# The Hot Jupiter Radius Anomaly and Stellar Connections

### v1

@@
Abstract
The extremely close proximity of hot Jupiters to their parent stars has dramatically affected both their atmospheres and interiors, inflating them to up to twice the radius of Jupiter. The physical mechanism responsible for this inflation remains unknown, though many proposals have been put forward. I will review the known hot Jupiter population, the proposed inflation mechanisms, and the evidence for and against them collected thus far. In doing so, I will cover the ways that hot Jupiter interiors may be simulated computationally in detail, and present some useful formulas for estimating their radii, heating, intrinsic temperature, and tentative magnetic field strength. I will also cover the related issues of hot Jupiter intrinsic temperatures and radiative-convective boundaries, the potential connection with planetary magnetic fields, and the effects of stellar tides on the planet. Finally, I conclude with the suggestion that more than one mechanism may be operating in concert with each other and propose various avenues for future progress in understanding these objects.
@@ 1 Introduction
The discovery of on close-in orbits around their parent stars (Mayor and Queloz 1995, the first being) came as something of a surprise and raised the question of how such objects could be formed (e.g., Pollack et al. 1996; Ida and Lin 2004). Some argued the planets could have formed in-situ (Bodenheimer et al. 2000, e.g.,), while others favored migration from wider separations due to interactions with the disk (e.g., Masset and Papaloizou 2003) or the star (eccentric migration Marzari and Weidenschilling 2002; Ford and Rasio 2006). This was the first of two great problems in physics. The second emerged when the transit method revealed a radius for HD 209458 b: (Charbonneau et al. 2000) and (Henry et al. 2000). This large size stood in stark contrast with our previous understanding of giant planet physics, which predicted a radius of (Bodenheimer et al. 2001; Guillot and Showman 2002).
Initially, this discrepancy was understood to be either the result of observational and theoretical uncertainties (Burrows et al. 2003, 2004) or as an exceptional case of a large planet (Bodenheimer et al. 2003) having an unusually hot interior due to tidal interactions with the parent star (Bodenheimer et al. 2001). Guillot and Showman (2002) proposed that while the planet indeed must be very hot (which was supported by Baraffe et al. 2003; Chabrier et al. 2004), the most likely source of heating was not tidal interactions but rather the intense instellation from the parent star heating the interior through a previously unknown physical mechanism. A small number of additional transiting hot Jupiters were discovered, including OGLE-TR-56 b (Konacki et al. 2003) and TrES-1 Alonso et al. (2004), revealing more moderately inflated hot Jupiters similar to HD 209458 b.
The number of known hot Jupiters grew substantially in the late 2000s thanks in large part to discoveries made by the Hungarian Automated Telescope (HAT; Bakos et al. 2004, 2007) and Wide Angle Search for Planets (WASP; Pollacco et al. 2006; Collier Cameron et al. 2007) surveys. These exhibited consistently larger radii than theoretical models could produce (Fig. 1), demonstrating that the radius anomaly was not a special case for the first few hot Jupiters, but instead a process broadly affecting the entire population. Subsequent discoveries using Kepler (Borucki et al. 2010), K2 (Howell et al. 2014), CoRoT (Auvergne et al. 2009), HAT-South (HATS; Bakos et al. 2013), the Transiting Exoplanet Survey Satellite (TESS; Ricker et al. 2014), and others have expanded the number of transiting hot Jupiters with mass measurements to over 500! In this chapter, I will give an overview of this population, explain how interior structure models of these objects are calculated, list some of the proposed explanations for their radii, and identify important observational implications of these hypotheses.
@@ 2 The Observed Population
The current set of giant exoplanets with known radii and masses is shown in Fig. 1, plotted as radius against flux. The black dashed line depicts the radius of a pure H/He object with no extra internal heating. As originally observed by Miller and Fortney (2011) and Demory and Seager (2011), radius inflation can be seen to turn on at erg s-1 cm-2 ( K), marked with a vertical dotted line. Above this radii are correlated with the incident flux. The most highly inflated planets reach at in the 1500-1800 K range, and are generally somewhat less massive than Jupiter. On average, more massive planets are smaller for a given incident flux.
@@
for planets with . is the planetary radius in , is the planetary mass, and is the incident flux onto the planet in erg s-1 cm-1.
For completeness and comparison we’ll also mention a few other empirical fits to hot Jupiter radii. Chen and Kipping (2017) provide a rigorously-derived mass-radius relation for a wide range of planetary masses and an easy-to-use code for calculating predictions in either direction, but note that they do not use the incident flux as a predictor. They therefore output radii corresponding with a typical flux of around erg s-1 cm-1. For hot Jupiters in particular, Thorngren et al. (2021) experimented with a large number of functional forms with a larger sample size but recommends a similar form to that of Weiss et al. (2013):
where is the radius in , is the mass in , and is the incident flux in erg s-1 cm-1. Note the addition of a cross-term between flux and radius: the inflationary effect of flux decreases with increasing mass.
In some cases it may be desirable to estimate the radius in the absence of anomalous heating – e.g. to calculate the radius anomaly as in Laughlin et al. (2011). Fortney et al. (2007) calculated a grid of model radii that may be interpolated for this purpose. Alternatively, (Thorngren et al. 2019b) estimate the following for giant planets with :
@@
Finally, we will mention a few hot Jupiters of particular interest. HD 209458 b was the first discovered hot Jupiter, but it also is a fairly typical example of the class that is extremely well-studied, making it a good choice if readers want a single ”typical case” to study. KELT-9 b (Gaudi et al. 2017) is the hottest known hot Jupiter with an impressive K, rivaling some K stars in surface temperature and thereby hosting an atmosphere with radically different chemistry from typical hot Jupiters (e.g., Hoeijmakers et al. 2018). The largest planet known is one of WASP-17 b, WASP-78 b, or WASP-79 b, all hot Jupiters of about 2 whose radius uncertainties prevent us from knowing which is largest. WASP-12 b, WASP-121 b, and WASP-103 b are believed to be the most tidally-distorted planets (Wahl et al. 2021), though by their radius vs Roche radius WASP-19 b and Qatar-10 b may be comparably distorted. Lastly, the shortest-period hot Jupiter with well-determined properties is currently TOI-2109 b (Wong et al. 2021), orbiting its star in just 16 hours.
2.1 Sampling Biases
@@
In this section, we have looked at the full sample of known transiting hot Jupiters, but one should bear in mind the observational biases that might be affecting it. Primarily, this refers to the biases of the transit method, whereby larger objects produce more easily detectable transit signals. Hot Jupiters are sufficiently large and short-period that for bright stars they are extremely likely to be successfully observed, limiting the importance of this bias. Indeed, Yee et al. (2021) found that for stars with , the (excluding the poles and galactic plane) is statistically nearly 100%! Further, they argue that TESS might reasonably expand this completeness to a limiting G-magnitude of about 12.5 (see also Sullivan et al. 2015; Zhou et al. 2021). This will present an exciting opportunity to better understand these planets, but we should still remember that the order in which they are discovered will not be random; the selection function for follow-up is undoubtedly complex, and unusual planets may be prioritized for confirmation.
@@
Another characteristic that biases the observed population of hot Jupiters is the subtle dependence on the age of the parent star. The principal cause of this is (Hamer and Schlaufman 2019), in which the closest-in hot Jupiters raise tides in their host stars that saps them of their angular momentum, decaying their orbits and eventually causing (Valsecchi et al. 2014, 2015). The tidal circularization of eccentric planets into hot Jupiters is expected to be generally very fast (Jackson et al. 2008), with a few rare cases of planets being observed presently undergoing this process, such as HD 80606 b (Naef et al. 2001). Finally, while higher-mass giant planets are thought to be fairly resistant to (Murray-Clay et al. 2009; Dwivedi et al. 2019), planets near Saturn’s mass or lower can still undergo substantial XUV-driven mass loss (Caldiroli et al. 2022; Kurokawa and Nakamoto 2014; Thorngren et al. 2023). As such, hot Jupiters orbiting young stars (e.g., Newton et al. 2019; Zhou et al. 2021) are a valuable point of comparison from before these processes have completed.
@@ 3 Modeling Hot Jupiter Interiors
The radius of a giant planet is mainly determined by the thermal state and composition of its interior, with only a small portion (, depending on the temperature) of the radius contributed by the atmosphere. As such, modeling these interiors is the first step in understanding the radius inflation problem. Even very modern computers are not fast enough to run fluid-dynamical simulations in three dimensions over billion-year timescales, so the long-term evolution of giant planets is relegated to one-dimensional models. In this section, we’ll review how one may construct such models of giant planets at a single point in time (static models), how these may be evolved forward through time, and additional relevant physics that one may wish to include in modeling these objects, depending on circumstance.
It should be mentioned that there are existing software packages that can be used to model , the most prominent being MESA (Paxton et al. 2013). While originally built for modeling the interiors and evolution of stars, the fundamental PDEs being solved are the same, allowing it to be successfully adapted to gas giant use cases on numerous occasions (Chen and Rogers 2016; Mankovich et al. 2016, e.g.,). Still, prospective users should be aware that these modifications are non-trivial. Another open-source option is MAGRATHEA (Huang et al. 2022); while mainly focused on the simulation of terrestrial or sub-Neptune planets, it can also be used for the simulation of true gas giant interiors.
@@ 3.1 Static Models
At an instant in time, the state of a giant planet can be described by its total mass and a complete description of the interior composition and temperature at a given mass shell . For a one-layer fully-convective planet, the composition can be the metal mass-fraction of the planet – 0 representing a metal-free planet and 1 being a planet composed entirely of metal (no hydrogen or helium). Thanks to our assumption of a fully-convective planet, the thermal state can be described by the specific entropy . An equation of state maps the specific entropy to a temperature for any given temperature, and the curve through temperature-pressure space is known as the adiabat. For giant planets, the of a hydrogen-helium envelope ranges from 5 (very “cold”) to 12 (very “hot”) per baryon, where is the Boltzmann constant. Giant planets begin their life with specific entropy somewhere in the 8-12 per baryon range (Marley et al. 2007; Cumming et al. 2018), which then decreases as energy is released through the atmosphere.
Given the mass, composition, and temperature structure of the planet, we can calculate its radius by solving the equations of hydrostatic equilibrium, mass conservation, and an appropriate (EOS):
Eq. 6 shows the across the equations of state of multiple components with mass fraction and density as a function of pressure and specific entropy. For these, one must decide what materials are to be modeled. A typical choice would start with hydrogen and helium at the solar (see Asplund et al. 2009) or protosolar ratios. For the latter, Serenelli and Basu (2010) give and , depending on the exact assumptions made. With a middling and (Asplund et al. 2009), we have . To obtain the density of H/He one should interpolate from an EOS table containing densities at various pressures and specific entropies. Chabrier et al. (1992) was the dominant choice for over a decade, but more recent ab-initio calculations (Nettelmann et al. 2008; Militzer and Hubbard 2013; Chabrier et al. 2019) have supplanted it. Chabrier and Debras (2021) would be a strong choice for H/He as of the time of writing, as it combines information from both Militzer and Hubbard (2013) and Chabrier et al. (2019). Fig. 3 shows an example interior model for a hot and warm Jupiter, calculated using these equations and the Chabrier et al. (2019) EOS.
@@
Multilayer Planets The discussion of the previous section has implicitly assumed that the planet in question is of uniform composition and convective throughout, but this is not necessarily the case. A major issue with this is that the core accretion model of planet formation requires a heavy-element core to form first in order to reach runaway gas accretion (Pollack et al. 1996). This has led to the use of a two layer model consisting of an isothermal metallic core and a convective H/He envelope (e.g., Marley et al. 2007; Lopez and Fortney 2014; Piso et al. 2015), especially in the sub-Neptune regime and for studies focused on the gas-accretion process (e.g., Lee et al. 2014). However, studies of the solar system planets have favored additional layers.
For Uranus and Neptune, these require at least the presence of an additional water-rich envelope layer between the core and outer envelope (see Fortney et al. 2010; Bailey and Stevenson 2021). Jupiter and Saturn are both believed to feature “fuzzy” cores (Wahl et al. 2017; Debras and Chabrier 2019; Movshovitz et al. 2020; Mankovich and Fuller 2021), which is to say that the composition gradually transitions from H/He poor to rich, rather than via a sharp boundary. This gradual transition may feature a staircase of , in which convective layers are separated by thin non-convecting boundaries with steep temperature and composition gradients (Leconte and Chabrier 2012). Such a staircase appears to be the natural outcome of initial composition gradients (Moll et al. 2016; Helled and Stevenson 2017; Fuentes and Cumming 2020), and could have a significant impact on the thermal evolution if the number of steps is large enough (Chabrier and Baraffe 2007; Leconte and Chabrier 2012). On the other hand, the number of layers evolves with time as layers exchange material and heat and potentially merge (Vazan et al. 2015, 2016), with the two diffusive fluxes across the boundary being positively related to one another (Moll et al. 2017). As such, even if planets formed with composition gradients that formed a large number of layers in a semi-convective staircase, the number of layers may have declined significantly as the planet subsequently evolved, reducing their impact on the present-day thermal structure of the planet.

### v3

@@
The Hot Jupiter Radius Anomaly and Stellar Connections
@@ 1 Introduction
The discovery of gas giant planets on close-in orbits around their parent stars (the first being 51 Pegasi b; Mayor and Queloz 1995) came as something of a surprise and raised the question of how such objects could be formed (e.g., Pollack et al. 1996; Ida and Lin 2004). Some argued the planets could have formed in-situ (Bodenheimer et al. 2000, e.g.,), while others favored migration from wider separations due to interactions with the disk (e.g., Masset and Papaloizou 2003) or the star (eccentric migration Marzari and Weidenschilling 2002; Ford and Rasio 2006). This was the first of two great problems in hot Jupiter physics. The second emerged when the transit method revealed a radius for HD 209458 b: (Charbonneau et al. 2000) and (Henry et al. 2000). This large size stood in stark contrast with our previous understanding of giant planet physics, which predicted a radius of (Bodenheimer et al. 2001; Guillot and Showman 2002).
Initially, this discrepancy was understood to be either the result of observational and theoretical uncertainties (Burrows et al. 2003, 2004) or as an exceptional case of a large planet (Bodenheimer et al. 2003) having an unusually hot interior due to tidal interactions with the parent star (Bodenheimer et al. 2001). Guillot and Showman (2002) proposed that while the planet indeed must be very hot (which was supported by Baraffe et al. 2003; Chabrier et al. 2004), the most likely source of heating was not tidal interactions but rather the intense instellation from the parent star heating the interior through a previously unknown physical mechanism. A small number of additional transiting hot Jupiters were discovered, including OGLE-TR-56 b (Konacki et al. 2003) and TrES-1 (Alonso et al. 2004), revealing more moderately inflated hot Jupiters similar to HD 209458 b.
The number of known hot Jupiters grew substantially in the late 2000s thanks in large part to discoveries made by the Hungarian Automated Telescope (HAT; Bakos et al. 2004, 2007) and Wide Angle Search for Planets (WASP; Pollacco et al. 2006; Collier Cameron et al. 2007) surveys. These exhibited consistently larger radii than theoretical models could produce (Fig. 1), demonstrating that the radius anomaly was not a special case for the first few hot Jupiters, but instead a process broadly affecting the entire population. Subsequent discoveries using Kepler (Borucki et al. 2010), K2 (Howell et al. 2014), CoRoT (Auvergne et al. 2009), HAT-South (HATS; Bakos et al. 2013), the Transiting Exoplanet Survey Satellite (TESS; Ricker et al. 2014), and others have expanded the number of transiting hot Jupiters with mass measurements to over 500! This chapter will give an overview of this population, explain how interior structure models of these objects are calculated, list some of the proposed explanations for their radii, and identify important observational implications of these hypotheses.
@@ 2 The Observed Population
The current set of giant exoplanets with known radii and masses is shown in Fig. 1, plotted as radius against flux. The black dashed line depicts the radius of a pure H/He object with no extra internal heating. As originally observed by Miller and Fortney (2011) and Demory and Seager (2011), radius inflation can be seen to turn on at erg s-1 cm-2 ( K), marked with a vertical dotted line. Above this radii are correlated with the incident flux. The most highly inflated planets reach at equilibrium temperatures in the 1500-1800 K range, and are generally somewhat less massive than Jupiter. On average, more massive planets are smaller for a given incident flux.
@@
for planets with . is the planetary radius in , is the planetary mass, and is the incident flux onto the planet in erg s-1 cm-2.
For completeness and comparison, a few other empirical fits to hot Jupiter radii bear mentioning. Chen and Kipping (2017) provide a rigorously-derived mass-radius relation for a wide range of planetary masses and an easy-to-use code for calculating predictions in either direction, but note that they do not use the incident flux as a predictor. They therefore output radii corresponding with a typical flux of around erg s-1 cm-2. For hot Jupiters in particular, Thorngren et al. (2021) experimented with a large number of functional forms with a larger sample size but recommends a similar form to that of Weiss et al. (2013):
where is the radius in , is the mass in , and is the incident flux in erg s-1 cm-2. Note the addition of a cross-term between flux and radius: the inflationary effect of flux decreases with increasing mass.
In some cases it may be desirable to estimate the radius in the absence of anomalous heating – e.g. to calculate the radius anomaly as in Laughlin et al. (2011). Fortney et al. (2007) calculated a grid of model radii that may be interpolated for this purpose. Alternatively, (Thorngren et al. 2019b) estimate the following mass-radius relation for giant planets with :
@@
Finally, it is worth noting a few hot Jupiters of particular interest. HD 209458 b was the first discovered hot Jupiter, but it also is a fairly typical example of the class that is extremely well-studied, making it a good choice if readers want a single ”typical case” to study. KELT-9 b (Gaudi et al. 2017) is the hottest known hot Jupiter with an impressive K, rivaling some K stars in surface temperature and thereby hosting an atmosphere with radically different chemistry from typical hot Jupiters (e.g., Hoeijmakers et al. 2018). The largest planet known is one of WASP-17 b, WASP-78 b, or WASP-79 b, all hot Jupiters of about 2 whose radius uncertainties prevent us from knowing which is largest. WASP-12 b, WASP-121 b, and WASP-103 b are believed to be the most tidally-distorted planets (Wahl et al. 2021), though by their radius vs Roche radius WASP-19 b and Qatar-10 b may be comparably distorted. Lastly, the shortest-period hot Jupiter with well-determined properties is currently TOI-2109 b (Wong et al. 2021), orbiting its star in just 16 hours.
2.1 Observational Biases
@@
This section has examined the full sample of known transiting hot Jupiters, but one should bear in mind the observational biases that might be affecting it. Primarily, this refers to the biases of the transit method, whereby larger objects produce more easily detectable transit signals. Hot Jupiters are sufficiently large and short-period that for bright stars they are extremely likely to be successfully observed, limiting the importance of this bias. Indeed, Yee et al. (2021) found that for stars with , the completeness (excluding the poles and galactic plane) is statistically nearly 100%! Further, they argue that TESS might reasonably expand this completeness to a limiting G-magnitude of about 12.5 (see also Sullivan et al. 2015; Zhou et al. 2021). This will present an exciting opportunity to better understand these planets, but one should still remember that the order in which they are discovered will not be random; the selection function for follow-up is undoubtedly complex, and unusual planets may be prioritized for confirmation. For planets that are observed, the large pixels of TESS (Ricker et al. 2014) could risk light curve contamination from background or companion stars (causing an underestimation of the planetary radius), but these cases can be caught and corrected for through either spectroscopy of the star or by comparison with the GAIA data (e.g. Mugrauer and Michel 2020).
@@
Another characteristic that biases the observed population of hot Jupiters is the subtle dependence on the age of the parent star. The principal cause of this is tidal inspiral (Hamer and Schlaufman 2019), in which the closest-in hot Jupiters raise tides in their host stars that saps them of their angular momentum, decaying their orbits and eventually causing Roche lobe overflow (Valsecchi et al. 2014, 2015). The tidal circularization of eccentric planets into hot Jupiters is expected to be generally very fast (Jackson et al. 2008), with a few rare cases of planets being observed presently undergoing this process, such as HD 80606 b (Naef et al. 2001). Finally, while higher-mass giant planets are thought to be fairly resistant to XUV-driven mass loss (Murray-Clay et al. 2009; Dwivedi et al. 2019), planets near Saturn’s mass or lower can still undergo substantial XUV-driven mass loss (Caldiroli et al. 2022; Kurokawa and Nakamoto 2014; Thorngren et al. 2023). As such, hot Jupiters orbiting young stars (e.g., Newton et al. 2019; Zhou et al. 2021) are a valuable point of comparison from before these processes have completed.
@@ 3 Modeling Hot Jupiter Interiors
The radius of a giant planet is mainly determined by the thermal state and composition of its interior, with only a small portion (, depending on the temperature) of the radius contributed by the atmosphere. As such, modeling these interiors is the first step in understanding the radius inflation problem. Even very modern computers are not fast enough to run fluid-dynamical simulations in three dimensions over billion-year timescales, so the long-term evolution of giant planets is relegated to one-dimensional models. This section will review how one may construct such models of giant planets at a single point in time (static models), how these may be evolved forward through time, and additional relevant physics that one may wish to include in modeling these objects, depending on circumstance.
It should be mentioned that there are existing software packages that can be used to model planetary interiors, the most prominent being MESA (Paxton et al. 2013). While originally built for modeling the interiors and evolution of stars, the fundamental PDEs (partial differential equations) being solved are the same, allowing it to be successfully adapted to gas giant use cases on numerous occasions (Chen and Rogers 2016; Mankovich et al. 2016, e.g.,). Still, prospective users should be aware that these modifications are non-trivial. Another open-source option is MAGRATHEA (Huang et al. 2022); while mainly focused on the simulation of terrestrial or sub-Neptune planets, it can also be used for the simulation of true gas giant interiors.
@@ 3.1 Static Models
At an instant in time, the state of a giant planet can be described by its total mass and a complete description of the interior composition and temperature at a given mass shell . For a one-layer fully-convective planet, the composition can be the metal mass-fraction of the planet : 0 representing a metal-free planet and 1 being a planet composed entirely of metal (no hydrogen or helium). Due to the assumption of a fully-convective planet, the thermal state can be described by the specific entropy . An equation of state maps the specific entropy to a temperature for any given pressure, and the curve through temperature-pressure space is known as the adiabat. For giant planets, the specific entropy of a hydrogen-helium envelope ranges from 5 (very “cold”) to 12 (very “hot”) per baryon, where is the Boltzmann constant. Giant planets begin their life with specific entropy somewhere in the 8-12 per baryon range (Marley et al. 2007; Cumming et al. 2018), which then decreases as energy is released through the atmosphere.
Given the mass, composition, and temperature structure of the planet, its radius can be calculated by solving the equations of hydrostatic equilibrium, mass conservation, and an appropriate equation of state (EOS):
Eq. 6 shows the additive volumes approximation across the equations of state of multiple components with mass fraction and density as a function of pressure and specific entropy. For these, one must decide what materials are to be modeled. A typical choice would start with hydrogen and helium at the solar (see Asplund et al. 2009) or protosolar ratios. For the latter, Serenelli and Basu (2010) give and , depending on the exact assumptions made. With a middling and (Asplund et al. 2009), we have . To obtain the density of H/He one should interpolate from an EOS table containing densities at various pressures and specific entropies. Chabrier et al. (1992) was the dominant choice for over a decade, but more recent ab-initio calculations (Nettelmann et al. 2008; Militzer and Hubbard 2013; Chabrier et al. 2019) have supplanted it. Chabrier and Debras (2021) would be a strong choice for H/He as of the time of writing, as it combines information from both Militzer and Hubbard (2013) and Chabrier et al. (2019).
@@
Multilayer Planets The discussion of the previous section has implicitly assumed that the planet in question is of uniform composition and convective throughout, but this is not necessarily the case. A major issue with this is that the core accretion model of planet formation requires a heavy-element core to form first in order to reach runaway gas accretion (Pollack et al. 1996). This has led to the use of a two layer model consisting of an isothermal metallic core and a convective H/He envelope (e.g., Marley et al. 2007; Lopez and Fortney 2014; Piso et al. 2015), especially in the sub-Neptune regime and for studies focused on the gas-accretion process (e.g., Lee et al. 2014). Fig. 3 shows an example interior model for a hot and warm Jupiter featuring an envelope surrounding a distinct rock-ice core.
For Uranus and Neptune, these require at least the presence of an additional water-rich envelope layer between the core and outer envelope (see Fortney and Nettelmann 2010; Bailey and Stevenson 2021). Jupiter and Saturn are both believed to feature “fuzzy” cores (Wahl et al. 2017; Debras and Chabrier 2019; Movshovitz et al. 2020; Mankovich and Fuller 2021), which is to say that the composition gradually transitions from H/He poor to rich, rather than via a sharp boundary. This gradual transition may feature a staircase of layered semiconvection, in which convective layers are separated by thin non-convecting boundaries with steep temperature and composition gradients (Leconte and Chabrier 2012). Such a staircase appears to be the natural outcome of initial composition gradients (Moll et al. 2016; Helled and Stevenson 2017; Fuentes et al. 2022), and could have a significant impact on the thermal evolution if the number of steps is large enough (Chabrier and Baraffe 2007; Leconte and Chabrier 2012). On the other hand, the number of layers evolves with time as layers exchange material and heat and potentially merge (Vazan et al. 2015, 2016), with the two diffusive fluxes across the boundary being positively related to one another (Moll et al. 2017). As such, even if planets formed with composition gradients that formed a large number of layers in a semi-convective staircase, the number of layers may have declined significantly as the planet subsequently evolved, reducing their impact on the present-day thermal structure of the planet.